1. Find the dimension of the subspace spanned by the following vectors in V3(R). (1, 0, 2), (2, 0, 1), (1, 0, 1).
1
Expert's answer
2021-10-06T13:33:27-0400

"Solution: 
\\We ~ know~ that~ Basis ~of ~vector ~space ~V ~is~ a ~linearly~ independent ~set~ that~ spans~ V.
\\ \therefore ~dimension~ of ~V = Card(basis ~of~ V)
\\Now ~to ~check ~ linearly ~independent~, we ~need ~to~ write~ given ~vectors~ in ~matrix ~
\\form ~and ~reduce~ it ~to ~row ~echelon ~form.
\\ \therefore ~Number ~of ~non- zero~ rows ~are~ the~ dimensions ~of~ the ~subspace~ spanned ~\\by ~the~ given~ vectors.
\\ \therefore~ We~ form~ a~ matrix ~whose ~rows~ are~ given ~by ~these ~vectors ~and~ reduce~ ~\\it ~to~ echelon~ form
\\ \therefore \begin{bmatrix}
1 & 0 & 2\\
2 & 0 & 1\\
1 & 0 & 1
\end{bmatrix}
\\Apply~ R_2 -2R_1 \rightarrow R_2
\\ ~~~~\begin{bmatrix}
1 & 0 & 2\\
0 & 0 & -3\\
1 & 0 & 1
\end{bmatrix}
\\ Apply ~R_3 -R_1 \rightarrow ~R_3 
\\ ~~~~\begin{bmatrix}
1 & 0 & 2\\
0 & 0 & -3\\
0 & 0 & -1
\end{bmatrix}
\\Apply~(R_3/-3) \rightarrow~R_3 
\\~~~~\begin{bmatrix}
1 & 0 & 2\\
0 & 0 & 1\\
0 & 0 & -1
\end{bmatrix}
\\Apply~R_3+R_2 \rightarrow R_3
\\~~~~\begin{bmatrix}
1 & 0 & 2\\
0 & 0 & 1\\
0 & 0 & 0
\end{bmatrix}
\\ Here~non-zero~ vectors ~are~2.
\\ \therefore 2 ~is ~the~ dimension~ of ~the~ subspace~ spanned~ by ~the~given ~vectors~ in V3(R)."